OpenStudy (anonymous):

Which value(s) must be excluded from solutions for this equation? 2/x^2-1 = 3/x+1 + 4/x-1

OpenStudy (anonymous):

-1 and +1 this is because the denominators can never equal 0, because we cannot divide by zero. the denominators here are x^1-1, x+1, x-1

OpenStudy (anonymous):

Well one of the answers is indeed 0, so why wouldn't I choose that?

OpenStudy (anonymous):

@robz8

OpenStudy (lopezking1):

2/ x^2-1 = 3/x+1 + 4/ x-1 2/(x^2-1) = (3x- 3 + 4x +4) / (x+1)(x-1) 2/(x^2-1) = (7x + 1) / (x^2 - 1) 2 = 7x +1 x= 1/7, (substitute the value, if the equation is not satisfied then x is not equal to 1/7)

OpenStudy (lopezking1):

@Cubi-Cal

OpenStudy (anonymous):

But I'm confused, the given answers are -1, 0, -1 and 1, or 1. @lopezking1

OpenStudy (lopezking1):

Ohh well then in this case it would be 0 i think

OpenStudy (anonymous):

Thanks!c:

OpenStudy (lopezking1):

Np